OpenStudy (anonymous):

Does 1/2 of 40 = 40 divided by 1/2?

OpenStudy (anonymous):

OH no

OpenStudy (anonymous):

1/2 of 40 =1/2 * 40 =20 40 divided by 1/2 =40/(1/2)=40*2=80

OpenStudy (anonymous):

no right

OpenStudy (anonymous):

Ya.... its no

OpenStudy (anonymous):

okay great just making sure :D